OpenStudy (anonymous):

On a set of house plans an architect represents a veranda of dimensions 7.2 m by 4.8 m by a rectangle 6 cm by 4 cm. What is the scale ratio used for the plans.

OpenStudy (anonymous):

l x b = k (x b) 7.2 x 4.8 = k (6 x 4) 34.56 = k (24) k = 34.56 /24 k = 1.44

OpenStudy (anonymous):

k is the scale factor

OpenStudy (anonymous):

the scale ratio is 1: 1.44